OpenStudy (pulsified333):

y=12F+Je^(t) where F and J are given constants. What is the derivative?

OpenStudy (kainashi):

The derivative of a constant =0 hence. \[y=12F + Je^f \] yields \[y'=12F\frac{ d }{ dt } + Je^t \frac{ d }{ dt } = 0+e^t\]

OpenStudy (kainashi):

sorry first e^f should be e^t Also I used chain rule onto the e^t \[e^t \frac{ d }{ dt } = e^t * t'\]

OpenStudy (pulsified333):

it isnt working :(

OpenStudy (pulsified333):

@satellite73

OpenStudy (kainui):

Kainashi is on the right track, but not quite there! Start with the function: \[y = 12F+Je^t\] Now take the derivative of both sides: \[\frac{dy}{dt} = \frac{d}{dt} ( 12 F +Je^t)\] The derivative has two important properties that we can use now: For any two functions, the derivative of the sum is the sum of the derivatives: \[\frac{d}{dt}( f(t)+g(t)) = \frac{d}{dt}( f(t)) + \frac{d}{dt}( g(t))\] And the other trick we have is that for any constant C we can pull it outside the derivative, and this is really just a consequence of the product rule since the derivative of a constant by itself is zero! \[\frac{d}{dt} ( C f(t)) = C \frac{d}{dt}(f(t))\] So try to use these tricks to solve your problem, give it your best shot.

OpenStudy (pulsified333):

the constant isnt C though its F and J

OpenStudy (kainui):

Yeah so now you can use these general rules to solve your problem by replacing C and f(t) with your functions.

OpenStudy (pulsified333):

FJ(12+e^t)?

OpenStudy (pulsified333):

I got it J(e^t)

OpenStudy (kainui):

Yeah that's the right answer

OpenStudy (kainui):

But what you just wrote right before that is not right, since \[FJ(12+e^t) \ne F12 + Je^t\]

OpenStudy (pulsified333):

yeah I know